OpenStudy (anonymous):

Find the indicated derivative using implicit differentiation. 4xy − y/4=1/x

OpenStudy (anonymous):

start with \[4y+4xy'-\frac{y'}{4}=-\frac{1}{x^2}\] then a bunch of algebra to solve for \(y'\)

OpenStudy (anonymous):

what do you mean by bunch?

OpenStudy (anonymous):

Group like terms, et al.

OpenStudy (anonymous):

well actually not that much subtract \(4y\) factor out the \(y'\) on the left and divide

OpenStudy (anonymous):

I finally got it \[-\frac{ 1 }{x ^{2} }-4y /4x-1/4\]

OpenStudy (anonymous):

Thanks!!